OpenStudy (anonymous):

Find all values of k that ensure that the given equation has exactly one solution. 4x2 + kx + 49 = 0 What is: k = ____ (smaller value) k = ____ (larger value)

OpenStudy (anonymous):

k^2 -4(4)(49) =0

OpenStudy (anonymous):

solve , you will get one positive and one negative , thus the largest and smallest

OpenStudy (anonymous):

k^2 - (4^2 )(7^2) = 0 k ^2 - 28^2 =0 k=+-28

OpenStudy (anonymous):

being a pro without a calculator :D

OpenStudy (anonymous):

make sure \[b^2-4ac=0\] with \[a=4,b=k,c=49\]

OpenStudy (anonymous):

\[k^2-4\times 4 \times 49=0\] \[k^2-4^2\times 7^2=0\] \[k^2=4^2\times 7^2\] \[k=\pm 4\times 7\] \[k= \pm 28\]

OpenStudy (anonymous):

what elecengineer wrote!